# Sigma: Semantically Informative Pre-training for Skeleton-based Sign Language Understanding (Under review)

### v1

@@ 3.2.3 Sign-grounded text matching and language modelling
To improve training efficiency and foster deeper cross-modal understanding, we propose an SGT encoder, inspired by Chen et al. (2020); Radford et al. (2021); Li et al. (2021; 2022a). It unifies sign-text matching (STM) and language modelling (LM), supporting dynamic alignment of visual and linguistic features within a single framework. A task-specific token guides the model to produce multimodal embeddings. A lightweight STM head, trained with binary cross-entropy, judges sign-text alignment. In parallel, the encoder autoregressively generates text with masked self-attention, with a cross-entropy LM loss enhancing language structure and semantics. To balance the synergy between matching and generation, we define a composite SGT loss:
@@ 4 Experiment
Datasets. We evaluate Sigma on a diverse set of benchmarks spanning different sign and spoken languages. WLASL2000 Li et al. (2020a) is used for ISLR evaluation, CSL-Daily Zhou et al. (2021a) serves as the benchmark both for CSLR and SLT. How2Sign Duarte et al. (2021) and OpenASL Shi et al. (2022) datasets are used for SLT evaluation (check Table 14 for their statistics and information).
@@ 4.3 Trade-off analysis between text-matching and language modelling
ST-GCN Yan et al. (2018) 34.40 32.53
HMA Hu et al. (2021b) 37.91 35.90
SignBERT Zhou et al. (2021c) 39.40 36.74
BEST Zhao et al. (2023) 46.25 43.52
SignBERT+ Hu et al. (2023a) 48.85 46.37
MSLU Zhou et al. (2024) 56.29 53.29
NLA-SLR Zuo et al. (2023) 61.05 58.05
Uni-Sign Li et al. (2025) 63.52 61.32
@@ Methods
SignBT Zhou et al. (2021a) 33.20 33.20
AdaBrowse Hu et al. (2023e) 31.20 30.70
SEN Hu et al. (2023d) 31.10 30.70
CorrNet Hu et al. (2023c) 30.60 30.10
MSLU Zhou et al. (2024) 28.60 27.90
CoSign Jiao et al. (2023) 28.10 27.20
Uni-Sign Li et al. (2025) 26.70 26.00
@@ 5 Comparison with state-of-the-art methods
We evaluate Sigma across the three aforementioned core SLU tasks. For ISLR on the WLASL2000 dataset, our model sets a new performance benchmark (see Table 5). These results demonstrate strong gesture recognition and effective feature discrimination. For CSLR on the CSL-Daily dataset, as shown in Table 7, Our method achieves new state-of-the-art (SOTA) performance, surpassing the strong pose-RGB-based Uni-Sign model, highlighting improved temporal modelling and more precise alignment between sign sequences and sign glosses relying solely on skeletal data. For SLT (see Table 9 and Table 9), Sigma shows strong performance across How2Sign, OpenASL, and CSL-Daily. On How2Sign, it delivers improvements across all evaluation metrics. Sigma achieves new SOTA results on OpenASL across all evaluation metrics used for this study. On CSL-Daily, it surpasses all gloss-free methods and rivals the long-standing gloss-based SOTA model CV-SLT. These results confirm the generalisability of Sigma across varied datasets and SLU tasks. The complete results are in Appendix I.
@@
GloFE-VN Lin et al. (2023) 14.90 2.20 12.60
YouTube-ASL Uthus et al. (2023) 37.80 12.40 -
MSLU Zhou et al. (2024) 20.10 2.40 17.20
SLT-IV Tarrés et al. (2023) 34.00 8.00 -
$C^{2}RL$ Chen et al. (2025) 29.10 9.40 27.00
FLa-LLM Chen et al. (2024) 29.80 9.70 27.80
@@
GloFE-VN Lin et al. (2023) 21.56 7.06 21.75
Conv-GRU Camgoz et al. (2018) 16.11 4.58 16.10
I3D-transformer Shi et al. (2022) 18.31 5.66 18.64
OpenASL Shi et al. (2022) 20.92 8.59 21.02
Uni-Sign Li et al. (2025) 49.35 23.14 43.22
$C^{2}RL$ Chen et al. (2025) 31.46 13.21 31.36
@@
SLRT Camgoz et al. (2020) 37.47 11.88 37.96 37.38 11.79 36.74
ConSLT Fu et al. (2023) - 14.80 41.46 - 14.53 40.98
SignBT Zhou et al. (2021a) 51.46 20.80 49.49 51.42 21.34 49.31
MMTLB Chen et al. (2022a) 53.81 24.42 53.38 53.31 23.92 53.25
SLTUNET Zhang et al. (2023) - 23.99 53.58 54.98 25.01 54.08
TS-SLT Chen et al. (2022b) 55.21 25.76 55.10 55.44 25.79 55.72
CV-SLT Zhao et al. (2024a) 56.36 28.24 56.36 58.29 28.94 57.06
@@
SLRT Camgoz et al. (2020) 21.03 4.04 20.51 20.00 3.03 19.67
GASLT Yin et al. (2023) - - - 19.90 4.07 20.35
MSLU Zhou et al. (2024) 33.28 10.27 33.13 33.97 11.42 33.80
NSLT Camgoz et al. (2018) 34.22 7.96 34.28 34.16 7.56 34.54
GFSLT-VLP Zhou et al. (2023) 39.20 11.07 36.70 39.37 11.00 36.44
FLa-LLM Chen et al. (2024) - - - 37.13 14.20 37.25
$C^{2}RL$ Chen et al. (2025) - - - 49.32 21.61 48.21
Uni-Sign Li et al. (2025) 55.30 26.25 56.03 55.08 26.36 56.51
SignLLM Gong et al. (2024) 42.45 12.23 39.18 39.55 15.75 39.91
Sign2GPT Wong et al. (2024) - - - 41.75 15.40 42.36
@@ 6 Conclusion
SLU requires the ability to recognise fine-grained visual patterns while simultaneously modelling complex linguistic semantics. We identify three key challenges in current SLU: 1) weak semantic grounding in visual features, 2) imbalanced local-global modelling, and 3) inadequate cross-modal alignment. To address these, we propose Sigma, a skeleton-based unified framework for semantically informative and transferable representation learning for SLU. Sigma introduces: 1) the SignEF mechanism for bidirectional visual-textual interaction during encoding, 2) the HAL strategy for optimising local and global alignment via contrastive objectives, and 3) a unified pre-training scheme combining contrastive learning, text matching, and language modelling. We validate the effectiveness of Sigma on multiple SLU benchmarks, including WLASL, CSL-Daily, How2Sign, and OpenASL. Sigma consistently exhibits strong performance across the aforementioned SLU tasks. These results underscore the importance of semantically informed pre-training for building scalable and robust skeleton-based SLU model.
@@ Appendix A Motivation
Appendix B Qualitative analysis
@@
For SLT, qualitative examples from English-based datasets (Tables12 and 11) demonstrate the ability of the model to produce fluent and contextually appropriate sentences and the generalization of the model across speaker identities as well as conversational styles. For instance in Table 11, it converts “I hope you’re having fun.” into “I hope you had fun with it.” showing an understanding of tense and implied context. These changes demonstrate that the model captures deeper semantic meaning rather than relying only on surface-level similarity. In Table 12, the model outputs “You will also talk with your legislators” instead of the reference “Also, be sure you talk to your legislators.” Though not a verbatim match, the generated sentence is syntactically sound and preserves the core message, demonstrating sentence-level comprehension. This reflects the impact of our HAL strategy in bridging local-global semantics across modalities. In Table 13, the Chinese examples show similar results. In the fourth row, “他醒来时发现自己在医院里” and “他醒来后,发现自己在医院” use different syntax but express the same idea. In the fifth example, “需要回家拿” and “回家要拿” describe the same action with a slight variation in sentence structure. In the final row, the model adds a rhetorical question “你想喝什么?” , which enhances the sentence without changing its meaning. These variations may affect token-based metrics, but they reflect natural language use and communicative clarity.
Appendix C Cluster aggregator
@@
To address the challenges of weak semantic grounding and limited cross-modal alignment, we design a cluster aggregator module inspired by Hou et al. (2024) to produce cluster-level textual embeddings that better correspond to visual sign units. Given a text input such as “curiosity,” the tokenizer splits it into subword tokens, which are then processed by the text encoder to generate token-level features. The aggregator groups these features into semantically coherent clusters. An offset calculator maps each original token to its cluster index, and the aggregation helper combines features within each cluster to form a compact representation. This process yields text embeddings that preserve semantic structure while reducing redundancy. Sigma supports both fine-grained gesture recognition and high-level translation. This mechanism contributes to more effective cross-modal learning and helps bridge the gap between dynamic visual input and structured language representations.
Appendix D Sign-grounded text encoder
@@
To mitigate the impact of weak semantic grounding and ineffective cross-modal alignment, we enhance text representations by integrating visual cues from sign features through a dual-path architecture. This SGT encoder consists of two parallel branches: a sign-text matching ($STM$) path and a language modelling ($LM$) path. The $STM$ path, repeated $M$ times, cooperates cross-attention layers where textual tokens attend to sign features, allowing the model to align linguistic units with visual semantics and enrich textual embeddings with SL gesture context. The $LM$ path, repeated $N$ times, uses standard transformer blocks with self-attention and feed-forward layers to preserve language fluency and syntactic structure. This dual-path setup enables the SGT encoder to learn representations that are both semantically grounded in visual input and linguistically coherent. During fine-tuning, all parameters except for the self-attention layers within the $STM$ path are transferred, ensuring effective knowledge reuse while allowing flexible adaptation to downstream SLU tasks. This design supports stronger cross-modal alignment and helps mitigate the semantic disconnect between dynamic sign inputs and static textual outputs.
Appendix E Skeletal data
@@
The sign sequences are skeletal data extracted using RTMPose Jiang et al. (2023) from MMPose Sengupta et al. (2020). Figure 8 illustrates the visualization of 69 keypoints per frame, including 21 for each hand, 9 for the body, and 18 for the face.
@@
Appendix F What makes annotations costly in sign language processing?
@@
Appendix G Rethinking the role of glosses in SheLU
@@
Gloss annotations have long been used as an intermediate representation in sign language translation, and they provide efficient and powerful supervision. By reducing the gap between raw visual input and spoken language output, glosses offer a structured, linguistically meaningful signal that has boosted SLT performance compared to purely end-to-end gloss-free approaches Zhou et al. (2023). At the same time, this benefit comes with significant drawbacks that increasingly limit scalability and linguistic fidelity. First, glosses are costly to obtain and difficult to scale. Producing them requires expert annotators and fine-grained temporal alignment, making data collection expensive and slow, and constraining the size of available datasets. Second, glosses act as an information bottleneck. A gloss sequence compresses rich and continuous sign expressions into discrete tokens, discarding nuances not represented in the gloss inventory and weakening the direct mapping from visual input to textual semantics. Third, gloss-based pipelines suffer from error propagation and mismatched objectives. Since they typically rely on a continuous sign language recognition front end, recognition errors are passed into the translation stage, while training remains split across recognition and translation tasks rather than being optimised jointly. Moreover, glosses often fail to capture divergences between sign language structure and the linear order of spoken languages. By committing to a single gloss sequence early, models risk locking in alignment hypotheses that hinder later reordering and discourse modelling. Another limitation lies in the inability of glosses to encode non-manual signals such as facial expressions, or the compositional use of multiple articulators, both of which carry critical linguistic meaning. In addition, gloss conventions vary across datasets and languages, introducing inconsistencies that reflect annotation practices rather than genuine linguistic differences, which in turn hinders transfer learning and cross-corpus pre-training. Finally, reliance on gloss labels restricts data efficiency. Although gloss supervision can enhance performance when available, it blocks the use of large unlabelled video corpora, whereas direct visual–text modelling allows learning from broader resources. Taken together, these issues highlight that glosses provide strong but narrow supervision: while effective in guiding alignment, they remain a bottleneck for scalability. This motivates our use of cluster-wise contrastive learning, which produces gloss-like groupings automatically and retains the advantages of structured alignment while avoiding the limitations of manual gloss annotation.
Appendix H Additional experiment
H.1 Optimising local cluster-wise contrastive learning strategies

### v2

@@ 3.2.3 Sign-grounded text matching and language modelling
To improve training efficiency and foster deeper cross-modal understanding, we propose an SGT encoder, inspired by (chen2020simple; radford2021learning; li2021align; li2022blip). It unifies sign-text matching (STM) and language modelling (LM), supporting dynamic alignment of visual and linguistic features within a single framework. A task-specific token guides the model to produce multimodal embeddings. A lightweight STM head, trained with binary cross-entropy, judges sign-text alignment. In parallel, the encoder autoregressively generates text with masked self-attention, with a cross-entropy LM loss enhancing language structure and semantics. To balance the synergy between matching and generation, we define a composite SGT loss:
@@ 4 Experiment
Datasets. We evaluate Sigma on a diverse set of benchmarks spanning different sign and spoken languages. WLASL2000 (li2020word) is used for ISLR evaluation, CSL-Daily (zhou2021improving) serves as the benchmark both for CSLR and SLT. How2Sign (duarte2021how2sign) and OpenASL (shi2022open) datasets are used for SLT evaluation (check Table 14 for their statistics and information).
@@ 4.3 Trade-off analysis between text-matching and language modelling
ST-GCN (yan2018spatial) 34.40 32.53
HMA (hu2021hand) 37.91 35.90
SignBERT (zhou2021signbert) 39.40 36.74
BEST (zhao2023best) 46.25 43.52
SignBERT+ (hu2023signbert+) 48.85 46.37
MSLU (zhou2024scaling) 56.29 53.29
NLA-SLR (zuo2023natural) 61.05 58.05
Uni-Sign (Li2025sign) 63.52 61.32
@@ Methods
SignBT (zhou2021improving) 33.20 33.20
AdaBrowse (hu2023adabrowse) 31.20 30.70
SEN (hu2023self) 31.10 30.70
CorrNet (hu2023continuous) 30.60 30.10
MSLU (zhou2024scaling) 28.60 27.90
CoSign (jiao2023cosign) 28.10 27.20
Uni-Sign (Li2025sign) 26.70 26.00
@@ 5 Comparison with state-of-the-art methods
We evaluate Sigma across the three aforementioned core SLU tasks. For ISLR on the WLASL2000 dataset, our model sets a new performance benchmark (see Table 5). These results demonstrate strong gesture recognition and effective feature discrimination. For CSLR on the CSL-Daily dataset, as shown in Table 7, Our method achieves new state-of-the-art (SOTA) performance, surpassing the strong pose-RGB-based Uni-Sign model, highlighting improved temporal modelling and more precise alignment between sign sequences and sign glosses relying solely on skeletal data. For SLT (see Table 9 and Table 9), Sigma shows strong performance across How2Sign, OpenASL, and CSL-Daily. On How2Sign, it delivers improvements across all evaluation metrics. Sigma achieves new SOTA results on OpenASL across all evaluation metrics used for this study. On CSL-Daily, it surpasses all gloss-free methods and rivals the long-standing gloss-based SOTA model CV-SLT. These results confirm the generalisability of Sigma across varied datasets and SLU tasks. The complete results are in Appendix J.
@@
GloFE-VN (lin2023gloss) 14.90 2.20 12.60
YouTube-ASL (uthus2023youtube) 37.80 12.40 -
MSLU (zhou2024scaling) 20.10 2.40 17.20
SLT-IV (tarres2023sign) 34.00 8.00 -
$C^{2}RL$ (chen2025c) 29.10 9.40 27.00
FLa-LLM (chen2024factorized) 29.80 9.70 27.80
@@
GloFE-VN (lin2023gloss) 21.56 7.06 21.75
Conv-GRU (camgoz2018neural) 16.11 4.58 16.10
I3D-transformer (shi2022open) 18.31 5.66 18.64
OpenASL (shi2022open) 20.92 8.59 21.02
Uni-Sign (Li2025sign) 49.35 23.14 43.22
$C^{2}RL$ (chen2025c) 31.46 13.21 31.36
@@
SLRT (camgoz2020sign) 37.47 11.88 37.96 37.38 11.79 36.74
ConSLT (fu2023token) - 14.80 41.46 - 14.53 40.98
SignBT (zhou2021improving) 51.46 20.80 49.49 51.42 21.34 49.31
MMTLB (chen2022simple) 53.81 24.42 53.38 53.31 23.92 53.25
SLTUNET (zhang2023sltunet) - 23.99 53.58 54.98 25.01 54.08
TS-SLT (chen2022two) 55.21 25.76 55.10 55.44 25.79 55.72
CV-SLT (zhao2024conditional) 56.36 28.24 56.36 58.29 28.94 57.06
@@
SLRT (camgoz2020sign) 21.03 4.04 20.51 20.00 3.03 19.67
GASLT (yin2023gloss) - - - 19.90 4.07 20.35
MSLU (zhou2024scaling) 33.28 10.27 33.13 33.97 11.42 33.80
NSLT (camgoz2018neural) 34.22 7.96 34.28 34.16 7.56 34.54
GFSLT-VLP (zhou2023gloss) 39.20 11.07 36.70 39.37 11.00 36.44
FLa-LLM (chen2024factorized) - - - 37.13 14.20 37.25
$C^{2}RL$ (chen2025c) - - - 49.32 21.61 48.21
Uni-Sign (Li2025sign) 55.30 26.25 56.03 55.08 26.36 56.51
SignLLM (gong2024llms) 42.45 12.23 39.18 39.55 15.75 39.91
Sign2GPT (wong2024sign2gpt) - - - 41.75 15.40 42.36
@@ 6 Conclusion
SLU requires the ability to recognise fine-grained visual patterns while simultaneously modelling complex linguistic semantics. We identify three key challenges in current SLU: 1) weak semantic grounding in visual features, 2) imbalanced local-global modelling, and 3) inadequate cross-modal alignment. To address these, we propose Sigma, a skeleton-based unified framework for semantically informative and transferable representation learning for SLU. Sigma introduces: 1) the SignEF mechanism for bidirectional visual-textual interaction during encoding, 2) the SignEF strategy for optimising local and global alignment via contrastive objectives, and 3) a unified pre-training scheme combining contrastive learning, text matching, and language modelling. We validate the effectiveness of Sigma on multiple SLU benchmarks, including WLASL, CSL-Daily, How2Sign, and OpenASL. Sigma consistently exhibits strong performance across the aforementioned SLU tasks. These results underscore the importance of semantically informed pre-training for building scalable and robust skeleton-based SLU model.
@@ Appendix A Motivation
Appendix B Advances in skeleton-based SLU
A growing line of work explores skeleton as a compact and semantically informative modality for SLU. Early work such as GCN-BERT (tunga2021pose) integrates graph convolutional networks over human joint graphs with transformer encoders, modelling spatial–temporal cues from skeletal sequences. Although effective for isolated SLR, it remains fully supervised and task-specific. To the best of our acknowledged, SignBERT hu2021signbert is the first work applies self-supervised pre-training for hand-centric skeletal representations. By masking and reconstructing hand trajectories and leveraging an explicit hand-shape model for regularisation, it learns richer visual embeddings and improves both isolated and continuous recognition. Building on stronger structural modelling, BEST (zhao2023best) advances skeleton-based pre-training by grouping body and hands into skeletal triplets and adopting a BERT-style masked unit modelling approach. A discrete VAE is used to tokenise continuous skeletal units into pseudo-tokens, enabling cross-entropy reconstruction and encouraging contextual reasoning over articulated hand–body interactions. BEST (zhao2023best) demonstrates strong generalisation across isolated SLR benchmarks. SignBERT+ hu2023signbertplus further extends this family of work by incorporating linguistic signals, and refining the pre-training tasks to support SLR and SLT. Compared with its predecessor, it provides more structured multi-task learning and better alignment between skeletal sequences and semantics. These methods reveal the strong potential of skeleton-only pre-training. However, they focus on capturing visual cues, lack joint modeling on the textual information, and remain visually grounded and largely task-specific. In contrast, Sigma builds on this foundation while moving toward a unified SLU paradigm and learns cross-modal alignment suitable for ISLR, CSLR, and SLT within a single framework.
Appendix C Qualitative analysis
@@
For SLT, qualitative examples from English-based datasets (Tables12 and 11) demonstrate the ability of the model to produce fluent and contextually appropriate sentences and the generalization of the model across speaker identities as well as conversational styles. For instance in Table 11, it converts “I hope you’re having fun.” into “I hope you had fun with it.” showing an understanding of tense and implied context. These changes demonstrate that the model captures deeper semantic meaning rather than relying only on surface-level similarity. In Table 12, the model outputs “You will also talk with your legislators” instead of the reference “Also, be sure you talk to your legislators.” Though not a verbatim match, the generated sentence is syntactically sound and preserves the core message, demonstrating sentence-level comprehension. This reflects the impact of our SignEF strategy in bridging local-global semantics across modalities. In Table 13, the Chinese examples show similar results. In the fourth row, “他醒来时发现自己在医院里” and “他醒来后,发现自己在医院” use different syntax but express the same idea. In the fifth example, “需要回家拿” and “回家要拿” describe the same action with a slight variation in sentence structure. In the final row, the model adds a rhetorical question “你想喝什么?” , which enhances the sentence without changing its meaning. These variations may affect token-based metrics, but they reflect natural language use and communicative clarity.
Appendix D Cluster aggregator
@@
To address the challenges of weak semantic grounding and limited cross-modal alignment, we design a cluster aggregator module inspired by (hou2024bagformer) to produce cluster-level textual embeddings that better correspond to visual sign units. Given a text input such as “curiosity,” the tokenizer splits it into subword tokens, which are then processed by the text encoder to generate token-level features. The aggregator groups these features into semantically coherent clusters. An offset calculator maps each original token to its cluster index, and the aggregation helper combines features within each cluster to form a compact representation. This process yields text embeddings that preserve semantic structure while reducing redundancy. Sigma supports both fine-grained gesture recognition and high-level translation. This mechanism contributes to more effective cross-modal learning and helps bridge the gap between dynamic visual input and structured language representations.
Appendix E Sign-grounded text encoder
@@
To mitigate the impact of weak semantic grounding and ineffective cross-modal alignment, we enhance text representations by integrating visual cues from sign features through a dual-path architecture. This SGT encoder consists of two parallel branches: a sign-text matching (STM) path and a language modelling (LM) path. The STM path, repeated $M$ times, cooperates with cross-attention layers where textual tokens attend to sign features, allowing the model to align linguistic units with visual semantics and enrich textual embeddings with SL gesture context. The LM path, repeated $N$ times, uses standard transformer blocks with self-attention and feed-forward layers to preserve language fluency and syntactic structure. This dual-path setup enables the SGT encoder to learn representations that are both semantically grounded in visual input and linguistically coherent. During fine-tuning, all parameters except for the self-attention layers within theSTM path are transferred, ensuring effective knowledge reuse while allowing flexible adaptation to downstream SLU tasks. This design supports stronger cross-modal alignment and helps mitigate the semantic disconnect between dynamic sign inputs and static textual outputs.
Appendix F Skeletal data
@@
The sign sequences are skeletal data extracted using RTMPose (jiang2023rtmpose) from MMPose (sengupta2020mm). Figure 8 illustrates the visualization of 69 keypoints per frame, including 21 for each hand, 9 for the body, and 18 for the face.
@@
Appendix G What makes annotations costly in sign language processing?
@@
Appendix H Rethinking the role of glosses in SLU
@@
Gloss annotations have long been used as an intermediate representation in sign language translation, and they provide efficient and powerful supervision. By reducing the gap between raw visual input and spoken language output, glosses offer a structured, linguistically meaningful signal that has boosted SLT performance compared to purely end-to-end gloss-free approaches (zhou2023gloss). At the same time, this benefit comes with significant drawbacks that increasingly limit scalability and linguistic fidelity. First, glosses are costly to obtain and difficult to scale. Producing them requires expert annotators and fine-grained temporal alignment, making data collection expensive and slow, and constraining the size of available datasets. Second, glosses act as an information bottleneck. A gloss sequence compresses rich and continuous sign expressions into discrete tokens, discarding nuances not represented in the gloss inventory and weakening the direct mapping from visual input to textual semantics. Third, gloss-based pipelines suffer from error propagation and mismatched objectives. Since they typically rely on a continuous sign language recognition front end, recognition errors are passed into the translation stage, while training remains split across recognition and translation tasks rather than being optimised jointly. Moreover, glosses often fail to capture divergences between sign language structure and the linear order of spoken languages. By committing to a single gloss sequence early, models risk locking in alignment hypotheses that hinder later reordering and discourse modelling. Another limitation lies in the inability of glosses to encode non-manual signals such as facial expressions, or the compositional use of multiple articulators, both of which carry critical linguistic meaning. In addition, gloss conventions vary across datasets and languages, introducing inconsistencies that reflect annotation practices rather than genuine linguistic differences, which in turn hinders transfer learning and cross-corpus pre-training. Finally, reliance on gloss labels restricts data efficiency. Although gloss supervision can enhance performance when available, it blocks the use of large unlabelled video corpora, whereas direct visual–text modelling allows learning from broader resources. Taken together, these issues highlight that glosses provide strong but narrow supervision: while effective in guiding alignment, they remain a bottleneck for scalability. This motivates our use of cluster-wise contrastive learning, which produces gloss-like groupings automatically and retains the advantages of structured alignment while avoiding the limitations of manual gloss annotation.
Appendix I Additional experiment
I.1 Optimising local cluster-wise contrastive learning strategies